# How to train your VAE

### v2

@@ 3.4 Network architecture
Fig. 3 depicts the components in our VAE implementation. The architecture of a VAE adheres to a foundational framework, depicted in Fig. 1. Within this framework, the two fundamental components of an AE persist: the Encoder (ENC) and the Decoder (DEC). However, the hallmark that sets the VAE apart is introducing a distinctive element: the sampler. We based on ResNetV2 [22] to implement our ResidualVAE. In particular, the Identity Residual Block (Id-ResBlock), Convolutional Residual Blocks (Conv-ResBlock) and Convolutional Residual Transposed Blocks (Conv-ResBlock Trans). The ”Id-ResBlock” maintains input and output dimensions, the ”Conv-ResBlock” reduces dimensions through a stride in the convolutional layer, and ”Conv-ResBlock Trans” extends the spatial dimensions through an interpolation performed by a Transposed-Convolution layer.
@@ 4 Experiments
We show experiments designed to demonstrate that our proposal preserves the essential characteristics of VAEs while generating high-quality data. That is, the generated image is similar to the image that generated the latent variable, small variations in the latent variables introduce slight variations in the generated faces and that the interpolation of latent vectors produces realistic data with smooth transitions.
@@
In the presented cases we use the Adam optimization method for trining the Residual-VAE and the Discriminator, we set the learning rate equal to $10^{-4}$. To demonstrate our proposal’s performance, we generate faces of $256\times 256$ pixels in RGB using the CelebA-HD database [23]. This database consists of 30,000 images that we split, after the alphanumeric ordering of files, in the first 24,000 for training and the reminder 6000 for testing.
@@
Our first experiment demonstrates the generalization capacity of our proposal. We generate variants of a random face $\bf x$$^{\prime}$, from the test set. We obtain the variations by sampling the posterior approximation: ${\mathbf{z}}^{\prime}\sim q_{\phi}({\bf z}|{{\mathbf{x}}=\bf x}^{\prime}).$ Next, we select two entries from the latent vector: $i,j\in[1,m]$. Then, with all inputs of $\mathbf{z}$$^{\prime}$ fixed except ${\mathbf{z}}_{i}^{\prime}$ and ${\mathbf{z}}_{j}^{\prime}$, we generate images that sample the probability ($\hat{\bf x}\sim p_{\theta}({{\mathbf{x}}|{\mathbf{z}}^{\prime}})$ varying these entries for combinations of ${\mathbf{z}}_{i}^{\prime}=[-20\,\sigma^{(i)},-10\,\sigma^{(i)},0,-10\,\sigma^{(i)},-20\,\sigma^{(i)}]$ and likewise for ${\mathbf{z}}_{j}^{\prime}$. Fig. 4 shows the generated variants. As we can see, the generated faces conserve a large part of the central face. In the illustrated example, the variables ${\mathbf{z}}_{0}$ and ${\mathbf{z}}_{2}$ correspond to the horizontal and vertical axes, respectively.
Fig. 4: VAE generated variants.
@@
Another interesting feature of VAEs is the ability to generate transitions between two faces, where the convex combination of two latent variables corresponds to a face with characteristics smoothly transitioning from one face to another. Fig. 5 depits generated transitions. Each row corresponds to the transition between two latent variables ${\bf z}_{1}$ and ${\bf z}_{2}$, corresponding to randomly selected images in the test dataset. The columns correspond to ${\bf z}_{k}^{\prime}=(1-\alpha_{k})\,{\bf z}_{1}+\alpha_{k}\,{\bf z}_{2}$, with $\alpha_{k}=\frac{k}{(n-1)}$; for $k=0,1,\ldots,5$.
Fig. 5: VAE generated transitions: $\hat{\mathbf{x}}\sim p_{\theta}({\mathbf{x}}|{\mathbf{z}}=\alpha\,{\mathbf{z}}^{(i)}+(1-\alpha)\,{\mathbf{z}}^{(j)})$, for $\alpha\in[0,1]$; see text.
@@
Finally, we show in Fig. 6 the joint histogram of two variables with typical estimated distributions. The histograms correspond to the latent variables ${\mathbf{z}}_{0}$ and ${\mathbf{z}}_{1}$ for 400 test data; on the right the marginal hisograms. It is notable that in our formulation 198 of 512 latent entries presented a bimodal histogram (with modes at -1 and 1). In the case of beta-VAE with beta less than 1, the collapsed histograms are monomodal and approach a Dirac delta; as Hinton noted. A more in-depth analysis of the use of these collapsed variables is beyond the scope of this article.
@@ 5 Conclusions
Our proposal substantially improves the reported results for standard VAE models while preserves the essential generative characteristics of the VAE. Although more complex variations of VAEs have been proposed (e.g., Hierarchical VAEs [5] and VQ-VAEs [8]), standard VAEs continue to be of great interest due to their elegant theoretical formulation and the promise of generating complex data from a simple distribution. Their main drawback is the posterior collapse, resulting in smoothed data. In this Work, we propose a solution to reduce these effects. Through experiments, we have demonstrated that reinterpreting the posterior as a mixture of Gaussians leads to a variant of the ELBO with a marginal computational cost. The list of our contributions in this work is as follows. First, an ELBO that reduces the collapse of the posterior towards the anterior (observed as the generation of very similar, blurry images). Our ELBO can be seen as an opposite version of Beta-VAE that preferentially weights quality reconstructions (with $\beta<1$), which is regularized to prevent individual posteriors from collapsing to a Dirac delta. Additionally, we use a VAE-PatchGAN training scheme that promotes realistic data through a discriminator that evaluates the realism of the generated data and reduces the contribution of the metric that compares pixel by pixel. Finally, we redesigned the architecture of the encoder and decoder to be of the Residual type to increase the encoding and decoding capabilities. Since the main contributions affect the training process, our results must be compared directly with those of a standard VAE or Beta-VAE. Qualitative comparisons of the results presented against those reported in the classic VAE formulations show superior performance of our proposal.
Fig. 6: Left, the join histogram of the latent variables $z_{0}$ and $z_{1}$ for the test dataset. On the right, the respective marginal histograms.
Acknowledges. Work supported in part by Conahcyt, Mexico (Grant CB-A1-43858).

### v3

@@ 3.4 Network architecture
Fig. 3 depicts the components in our VAE implementation. The architecture of a VAE adheres to a foundational framework, depicted in Fig. 1. Within this framework, the two fundamental components of an AE persist the Encoder (ENC) and the Decoder (DEC). However, the hallmark that sets the VAE apart is introducing a distinctive element: the sampler. We based our ResidualVAE implementation on ResNetV2 [21]. In particular, the Identity Residual Block (Id-ResBlock), Convolutional Residual Blocks (Conv-ResBlock), and Convolutional Residual Transposed Blocks (Conv-ResBlock-T). The ”Id-ResBlock” maintains input and output dimensions, the ”Conv-ResBlock” reduces dimensions through a stride in the convolutional layer, and ”Conv-ResBlock-T” extends the spatial dimensions through an interpolation performed by a transposed convolution layer.
@@ 4 Experiments
We show experiments designed to demonstrate that our proposal preserves the essential characteristics of VAEs while generating high-quality data. That is, the generated image is similar to the image that generated the latent variable; small variations in the latent variables introduce slight variations in the generated faces, and the interpolation of latent vectors produces realistic data with smooth transitions.
@@
In the presented cases we use the Adam optimization method for training the Residual-VAE and the Discriminator, we set the initial learning rate equal to $10^{-4}$. To demonstrate our proposal’s performance, we generate faces of $256\times 256$ pixels in RGB using the CelebA-HD database [22]. After the alphanumeric ordering of files, this database consists of 30,000 images that we split in the first 24,000 for training and the reminder 6000 for testing.
@@
Our first experiment demonstrates the generalization capacity of our proposal. We generate variants of a random face ${\bf x}^{\prime}$, from the test set. We obtain the variations by sampling the posterior approximation: ${\mathbf{z}}^{\prime}\sim q_{\phi}({\bf z}|{{\mathbf{x}}=\bf x}^{\prime}).$ Next, we select two entries from the latent vector: $a,b\in[1,m]$. Then, with all inputs of ${\mathbf{z}}^{\prime}$ fixed except ${\mathbf{z}}_{a}^{\prime}$ and ${\mathbf{z}}_{b}^{\prime}$, we generate images that sample the probability ($\hat{\bf x}\sim p_{\theta}({{\mathbf{x}}|{\mathbf{z}}^{\prime}})$ varying these entries for combinations of ${\mathbf{z}}_{a}^{\prime}=\mu_{a}^{(i)}+\delta_{k}\sigma_{a}^{(i)}$ with $\delta=[-20,0,20]$ and likewise for ${\mathbf{z}}_{b}^{\prime}$. Fig. 4 shows the generated variants. As we can see, the generated faces conserve a large part of the central face. In the illustrated example, the variables ${\mathbf{z}}_{0}$ and ${\mathbf{z}}_{2}$ correspond to the horizontal and vertical axes, respectively.
Fig. 4: Generated variants with the VAE using the proposed training scheme. Varying $z_{0}$ and $z_{1}$ according $\delta_{k}\times\sigma^{(i)}$ for $\delta=[-20,0,20]$.
@@
Another interesting feature of VAEs is the ability to generate transitions between two faces, where the convex combination of two latent variables corresponds to a face with characteristics smoothly transitioning from one face to another. Fig. 5 depicts generated transitions. Each row corresponds to the transition between two latent variables ${\bf z}_{1}$ and ${\bf z}_{2}$, corresponding to randomly selected images in the test dataset. The columns correspond to ${\bf z}_{k}^{\prime}=(1-\alpha_{k})\,{\bf z}_{1}+\alpha_{k}\,{\bf z}_{2}$, with $\alpha_{k}=\frac{k}{(n-1)}$; for $k=0,1,\ldots,5$.
Fig. 5: Transitions generated with VAEs by convex combinations of latent variables: $\hat{\mathbf{x}}\sim p_{\theta}({\mathbf{x}}|{\mathbf{z}}=\alpha\,{\mathbf{z}}^{(i)}+(1-\alpha)\,{\mathbf{z}}^{(j)})$, for $\alpha\in[0,1]$. First row, VAE trained with the proposed scheme. Second row, the same VAE model trained with the standard training strategy: $\beta=[1,5000]$ in (3)-(4), and $L_{1}$ norm for the negative log-likelihood.
@@
Finally, we show in Fig. 6 the joint histogram of two variables with typical estimated distributions. The histograms correspond to the latent variables ${\mathbf{z}}_{0}$ and ${\mathbf{z}}_{1}$ for 400 test data points; on the right the marginal histograms. It is notable that in our formulation, 198 of 512 latent entries presented a bimodal histogram (with modes at -1 and 1). In the case of beta-VAE with $\beta_{1}<\beta_{2}$, the collapsed histograms are monomodal and approach a Dirac delta, as noted in Ref. [15]. A more in-depth analysis of the use of these collapsed variables is beyond the scope of this article.
@@ 5 Conclusions
Our proposal substantially improves the reported results for standard VAE models while preserves the essential generative characteristics of the VAE. Although more complex variations of VAEs have been proposed (e.g., Hierarchical VAEs [3] and VQ-VAEs [6]), standard VAEs continue to be of great interest due to their elegant theoretical formulation and the promise of generating complex data from a simple distribution. Their main drawback is the posterior collapse problem, resulting in smoothed data. In this work, we propose a solution to reduce these effects. Through experiments, we have demonstrated that reinterpreting the posterior as a mixture of Gaussians leads to a variant of the ELBO with a marginal computational cost. The list of our contributions in this work is as follows. First, an ELBO that reduces the collapse of the posterior towards the anterior; observed as the generation of very similar blurry images. Our ELBO can be seen as an opposite version of Beta-VAE that preferentially weights quality reconstructions (with $\beta<1$), which is regularized to prevent individual posteriors from collapsing to a Dirac delta. Additionally, we use a VAE-PatchGAN training scheme that promotes realistic data through a discriminator that evaluates the generated data’s realism and reduces the metric’s contribution that compares pixel by pixel. Finally, we redesigned the encoder and decoder architecture to be of the Residual type to increase the encoding and decoding capabilities. Since the main contributions affect the training process, our results must be compared directly with those of a standard VAE or Beta-VAE. Qualitative comparisons of the results presented against those reported in the classic VAE formulations show the superior performance of our proposal. The code that implements the proposal is available in https://github.com/marianorivera/How2TrainUrVAE.
Fig. 6: Left, the join histogram of the latent variables $z_{0}$ and $z_{1}$ for the test dataset. On the right are the respective marginal histograms.
Acknowledges. Work supported in part by CONAHCYT, Mexico (Grant CB-A1-43858). The author thanks to the anonymous reviewers and L. M. Nunez for their useful comments to improve the quality of this paper.